OpenStudy (anonymous):

(x^2-2x+7)-(17x+x^2-8x)=3x-43 (5-4a+6a^2)-(4a-9a^2-5) Can u PLZ show me how to do these I have a test tommorow on polynomials.

myininaya (myininaya):

the first one is not right

OpenStudy (anonymous):

okay. do you know the distributive property. for example, what is 4(6) + 5(6)?

myininaya (myininaya):

\[(x^2-2x+7)-(17x+x^2-8x)=(x^2-2x+7)-(9x+x^2)=-2x-9x+7=-7x+7\]